# Towards Human-AI Complementarity with Predictions Sets

### v1

@@ 1 Introduction
In this context, Straitouri et al. [18, 19] have recently argued, both theoretically and empirically, that an alternative type of decision support systems may achieve human-AI complementarity, by design. Rather than providing a single label prediction and letting a human expert decide when and how to use the predicted label to update their own prediction, these systems provide a set of label predictions, namely a prediction set, and ask the expert to predict a label value from the set. [^2] There are many decision support systems used by experts that, under normal operation, forcefully limit experts’ level of agency. For example, in aviation, there are automated, adaptive systems that prevent pilots from taking certain actions based on the monitoring of the environment. To construct each prediction set, these systems rely on a conformal predictor [20, 21]. The conformal predictor first computes a non-conformity score for each potential label value using the output provided by a classifier (e.g., the softmax scores), and then adds a label value to the prediction set if its non-conformity score is below a data-driven threshold. Further, to optimize the performance of these systems, Straitouri et al. have introduced several methods to efficiently find the optimal value of the threshold used by the conformal predictor. However, it is unclear whether the optimal prediction sets under which human experts achieve the highest accuracy can always be constructed using a deterministic threshold rule as the one used by a conformal predictor. Motivated by this observation, in this work, our goal is to understand how to construct optimal prediction sets under which human experts achieve the highest accuracy.
@@
Our contributions. We first demonstrate that there exist (many) data distributions for which the optimal prediction sets under which the human experts achieve the highest accuracy cannot be constructed using a conformal predictor. Then, we show that the problem of finding the optimal prediction sets is $\mathsf{NP}$-hard by using a reduction from the $k$-clique problem [22]. More strongly, unless $\mathsf{P}=\mathsf{NP}$, we show that the problem is hard to approximate to any factor less than the size of the label set. However, we introduce a simple and efficient greedy algorithm that, for a large class of non-conformity scores and expert models, is guaranteed to find prediction sets that provably offer equal or greater performance than those constructed using conformal prediction. Moreover, using a simulation study with both synthetic and real expert predictions, we demonstrate that, in practice, our greedy algorithm finds near-optimal prediction sets offering greater performance than conformal prediction. [^3] We provide the code and data used in our experiments at the following link https://github.com/Networks-Learning/towards-human-ai-complementarity-predictions-sets under a liberal license.
@@ 2 Decision Support Systems Based on Prediction Sets
Given a multiclass classification task where, for each task instance, a human expert needs to predict the value of a ground truth label $y\in\mathcal{Y}=\{1,\ldots,L\}$, we focus on the design of a decision support system that, given a set of features $x\in\mathcal{X}$, helps the expert by narrowing down the set of potential label values to a subset of them ${\mathcal{S}}(x)\subseteq\mathcal{Y}$. Here, similarly as in Straitouri et al. [18, 19], we assume that, for any instance with features $x\in\mathcal{X}$, the system asks the expert’s prediction $\hat{y}\in\mathcal{Y}$ to belong to the prediction set ${\mathcal{S}}(x)$, i.e., $\hat{y}\in{\mathcal{S}}(x)$. The key rationale for restricting expert’s agency is that, if we would allow the expert to predict label values from outside the prediction set, good performance would depend on the expert developing a good sense of when to predict a label from the prediction set. Refer to Figure 1 for an illustration of the decision support system.
@@
Then, for any $x\in\mathcal{X}$, our goal is to find the optimal prediction set ${\mathcal{S}}^{*}(x)$ that maximizes the accuracy of the expert’s prediction, [^4] We denote random variables with capital letters and realizations of random variables with lowercase letters. i.e.,
@@ 3 On the Suboptimality of Conformal Prediction
Motivated by the above example, one may think of closing the above performance gap by incorporating information about the distribution of experts’ predictions in the definition of the non-conformity score. However, we cannot expect to fully close the performance gap since, as we will show next, the problem of finding the optimal prediction sets is NP-hard to solve and approximate to any factor less than the size of the label set $\mathcal{Y}$.
@@ 4       ${\mathcal{S}}_{k}\leftarrow\varnothing$
5       while $|{\mathcal{S}}_{k}|<k$ do
@@ 6             $\Delta^{*}\leftarrow-\infty$
8                   $\Delta\leftarrow\hat{g}({\mathcal{S}}_{k}\cup\{y\}\,|\,x)-\hat{g}({\mathcal{S}}_{k}\,|\,x)$
@@ 10                         $\Delta^{*}\leftarrow\Delta$, $y^{*}\leftarrow y$
12
13            ${\mathcal{S}}_{k}\leftarrow{\mathcal{S}}_{k}\cup\{y^{*}\}$
@@
15                   ${\mathcal{S}}\leftarrow{\mathcal{S}}_{k}$
16
17
18
19return ${\mathcal{S}}$
@@ 5 Experiments with Synthetic Data
Table 1: Empirical average test accuracy achieved by four different (simulated) human experts, each with a different noise value $\gamma$, on their own (None) and using prediction sets constructed with conformal prediction (Naive, Aps) and with the greedy algorithm (Greedy) on four synthetic classification tasks. In each classification task, the classifier $f$ used by conformal prediction and the greedy algorithm achieves a different average accuracy $P(Y^{\prime}=Y)$. The number of labels $L=10$, the size of the calibration set is $m=1000$, and we do not include brute force search because it achieves the same performance as the greedy algorithm. Each cell shows the average and standard deviation over $10$ runs. We denote the best results for each classification task in bold.
@@
Experimental setup. We create several synthetic multiclass classification tasks, each with $n=20$ features per sample and varying difficulty. Refer to Appendix C for more details about the classification tasks. For each classification task, we generate $18{,}000$ samples, which we split into a training set ($16{,}000$ samples), a calibration set ($1000$ samples), and a test set ($1000$ samples).
@@
Further, we use the calibration set to calibrate the (softmax) outputs of the logistic regression model $f$ using top-$k$-label calibration [52] with $k=5$. We also use it to estimate the confusion matrices $\bm{C}(\gamma)$ that parameterize the mixture of MNLs used to model (the predictions made by) the simulated human expert, and calculate the quantile $\hat{q}_{\alpha}$ used by conformal prediction. Finally, we use the test set to evaluate the average accuracy achieved by the simulated expert using prediction sets constructed with our greedy algorithm, brute force search and conformal prediction. Here, note that our greedy algorithm and brute force search have access to the true mixtures of MNLs used to model the simulated human expert. In our experiments, we implement conformal prediction using two popular non-conformity scores:
@@
For each non-conformity score and classification task, we report the results for the $\alpha$ value under which the expert achieves the highest average test accuracy and, to avoid empty sets, we always include the label value with the lowest non-conformity score in the prediction sets. Finally, we repeat each experiment ten times and, each time, we sample different training, calibration, and test sets.
Figure 2: (Left) Confusion matrix $\bm{C}$ for the predictions made by a (simulated) human expert on their own. The label $\bar{y}=\mathop{\mathrm{argmax}}_{y^{\prime}\neq y}C_{y^{\prime}y}$ that is most frequently mistaken with the ground truth-label $y$ is highlighted in red for $y\in\left\{0,2,6,8\right\}$. (Right) Empirical conditional probability that a prediction set includes $\{y,\bar{y}\}$ given $Y=y$ with conformal prediction (Naive and Aps) and our greedy algorithm (Greedy). In both panels, $\gamma=0.7$ and $\mathbb{P}(Y^{\prime}=Y)=0.7$.
@@
Results. We first estimate the average test accuracy achieved by four different (simulated) human experts, each with a different $\gamma$ value, on four classification tasks where the classifier $f$ achieves a different average accuracy $P(Y^{\prime}=Y)$. We report their average test accuracy on their own (None) and when using prediction sets constructed with conformal prediction (Naive, Aps), our greedy algorithm (Greedy) and brute force search (Brute Force Search). Table 1 summarizes the results, where we have not included brute force search because it achieves the same performance as our greedy algorithm. The results show that, using the greedy algorithm to construct prediction sets, the experts consistently achieve the highest average accuracy across classification tasks. Moreover, the results also show that, under the prediction sets constructed using the greedy algorithm, the average accuracy achieved by the expert degrades gracefully as $\gamma$ increases whereas, under the prediction sets constructed using conformal prediction, the average accuracy degrades significantly. Appendix D provides additional results for $L\in\{25,50\}$ suggesting that the relative gain in average accuracy provided by the greedy algorithm also increases with the number of labels and noise $\gamma$.
@@
To better understand why the prediction sets constructed by the greedy algorithm help human experts achieve higher average accuracy than those constructed conformal prediction, we now look closer into the structure of the prediction sets. Given a ground truth-label $Y=y$, let $\bar{y}=\mathop{\mathrm{argmax}}_{y^{\prime}\neq y}C_{y^{\prime}y}$ be the label that is most frequently mistaken with $y$. Then, we estimate the empirical conditional probability that a prediction set includes $\{y,\bar{y}\}$ given $Y=y$ with the greedy algorithm and conformal prediction. Figure 2 summarizes the results for $\gamma=0.7$ and $P(Y^{\prime}=Y)=0.7$ and $y\in\left\{0,2,6,8\right\}$. Appendix E include additional results for other configurations. The results show that, with the greedy algorithm, the empirical probability that a prediction set includes $\{y,\bar{y}\}$ given $Y=y$ is much lower (i.e., $2$-$3$x lower) than with conformal prediction despite it creates overall larger prediction sets.
@@ 6 Experiments with Real Data
In this section, we compare the average accuracy achieved by a simulated human expert using prediction sets constructed with our greedy algorithm, brute force search and conformal prediction on a real multiclass classification task over noisy natural images. The simulated human expert follows the mixture of MNLs introduced in Eq. 4, which is parameterized by the (estimated) confusion matrix of the predictions made by real human experts on their own. [^11] In Appendix F, we evaluate the goodness of fit of the mixture of MNLs to predictions made by real human experts using a support system based on prediction sets [19].
@@
Experimental setup. We experiment with the ImageNet16H dataset [7], which was created using $1{,}200$ natural images from the ImageNet Large Scale Visual Recognition Challenge (ILSRVR) 2012 dataset [54]. More specifically, in the ImageNet16H dataset, each of the above images was used to create four noisy images with different level of phase noise distortion $\omega\in\{80,95,110,125\}$ and the same ground-truth label $y$ from a label set $\mathcal{Y}$ of size $n=16$. In addition, for each noisy image, the dataset contains (approximately) six label predictions made by human experts on their own. In our experiments, we run and evaluate each method separately by grouping the above noisy images (and expert predictions) according to their level of noise. For each group of images and method, we use the deep neural network classifier VGG-19 [55] after $10$ epochs of fine-tuning as provided by Steyvers et al. [7]. Further, we randomly split the images (and expert predictions) in each group into two disjoint subsets, a calibration set ($800$ images), and a test set ($400$ images). We use the calibration set to (i) calibrate the (softmax) outputs of the VGG-19 scores using top-k-label calibration with $k=5$, (ii) estimate the confusion matrix $\bf C$ that parameterizes the mixture of MNLs used to model the simulated human expert, and (iii) calculate the quantile $\hat{q}_{\alpha}$ used by conformal prediction. We use the test set to evaluate the average accuracy the simulated expert achieves using prediction sets constructed with our greedy algorithm, brute force search and conformal prediction. Here, note that, similarly to in the experiments in synthetic data, our greedy algorithm and brute force search have access to the true mixture of MNLs used to model the simulated human expert. We implement conformal prediction using the same non-conformity scores used in the experiments with synthetic data and, for each non-conformity score and group of images, we report the results for the $\alpha$ value under which the expert achieves the highest average test accuracy. In Appendix G, we report results for all $\alpha$ values. To obtain error bars, we repeat each experiment $10$ times, sampling different calibration and test sets.
@@
Table 2: Average test accuracy achieved by a (simulated) human expert using the prediction sets constructed with conformal prediction (Naive and Aps), our greedy algorithm (Greedy), and brute force search (Brute Force Search) on the ImageNet16H dataset. Each cell shows the average and standard deviation over $10$ runs. We denote the best results in bold.

### v2

# Towards Human-AI Complementarity with Prediction Sets

@@ 1 Introduction
In this context, Straitouri et al. [18, 19] have recently argued, both theoretically and empirically, that an alternative type of decision support systems may achieve human-AI complementarity, by design. Rather than providing a single label prediction and letting a human expert decide when and how to use the predicted label to update their own prediction, these systems provide a set of label predictions, namely a prediction set, and ask the expert to predict a label value from the set. [^2] There are many decision support systems used by experts that, under normal operation, forcefully limit experts’ level of agency. For example, in aviation, there are automated, adaptive systems that prevent pilots from taking certain actions based on the monitoring of the environment. To construct each prediction set, these systems rely on a conformal predictor [20, 21]. The conformal predictor first computes a non-conformity score for each potential label value using the output provided by a classifier (e.g., the softmax scores), and then adds a label value to the prediction set if its non-conformity score is below a data-driven threshold computed using a calibration set. Further, to optimize the performance of these systems, Straitouri et al. have introduced several methods to efficiently find the optimal value of the threshold used by the conformal predictor. [^3] A threshold value is optimal if it maximizes the average accuracy achieved by an expert who predicts label values from the prediction sets created by the conformal predictor. However, it is unclear whether the optimal prediction sets maximizing the average accuracy achieved by an expert who uses such systems can always be constructed using a deterministic threshold rule as the one used by a conformal predictor. Motivated by this observation, in this work, our goal is to understand how to construct optimal prediction sets under which human experts achieve the highest average accuracy.
@@
Our contributions. We first demonstrate that there exist (many) data distributions for which the optimal prediction sets under which the human experts achieve the highest average accuracy cannot be constructed using a conformal predictor. Then, we show that the problem of finding the optimal prediction sets is $\mathsf{NP}$-hard by using a reduction from the $k$-clique problem [22]. More strongly, unless $\mathsf{P}=\mathsf{NP}$, we show that the problem is hard to approximate to any factor less than the size of the label set. However, we introduce a simple and computationally efficient greedy algorithm that, for a large class of non-conformity scores and expert models parameterized by a mixture of multinomial logit models (MNLs), is guaranteed to find prediction sets that provably offer equal or greater performance than those constructed using conformal prediction. Moreover, using a simulation study with both synthetic and real expert predictions, we demonstrate that, in practice, our greedy algorithm finds near-optimal prediction sets offering greater performance than conformal prediction. We have released an open-source implementation of our greedy algorithm as well as the code and data used in our experiments at https://github.com/Networks-Learning/towards-human-ai-complementarity-predictions-sets.
@@ 2 Decision Support Systems Based on Prediction Sets
Given a multiclass classification task where, for each task instance, a human expert needs to predict the value of a ground truth label $y\in\mathcal{Y}=\{1,\ldots,L\}$, we focus on the design of a decision support system that, given a set of features $x\in\mathcal{X}$, helps the expert by narrowing down the set of potential label values to a subset of them ${\mathcal{S}}(x)\subseteq\mathcal{Y}$. Here, similarly as in Straitouri et al. [18, 19], we assume that, for any instance with features $x\in\mathcal{X}$, the system asks the expert’s prediction $\hat{y}\in\mathcal{Y}$ to belong to the prediction set ${\mathcal{S}}(x)$, i.e., $\hat{y}\in{\mathcal{S}}(x)$. The key rationale for restricting the expert’s agency is that, if we would allow the expert to predict label values from outside the prediction set, a good performance would depend on the expert developing a good understanding of when to predict a label from the prediction set. In this context, it is worth highlighting that Straitouri et al. [19] run a large-scale human subject study to compare the above setting against an alternative setting where experts are allowed to predict label values from outside the prediction sets. They found that, in the alternative setting, the number of predictions in which the prediction sets do not contain the true label and the experts succeed is consistently smaller than the number of predictions in which the prediction sets contain the true label and the experts fail. As a consequence, in the alternative setting, experts perform worse. Refer to Figure 1 for an illustration of the decision support system.
@@
Then, for any $x\in\mathcal{X}$, our goal is to find the optimal prediction set ${\mathcal{S}}^{*}(x)$ that maximizes the average accuracy of the expert’s prediction, [^4] We denote random variables with capital letters and realizations of random variables with lowercase letters. i.e.,
@@ 3 On the Suboptimality of Conformal Prediction
Motivated by the above example, one may think of closing the above performance gap by incorporating information about the distribution of experts’ predictions in the definition of the non-conformity score. However, we cannot expect to fully close the performance gap since, as we will show next, the problem of finding the optimal prediction sets is $\mathsf{NP}$-hard to solve and approximate to any factor less than the size of the label set $\mathcal{Y}$.
@@ 4       ${\mathcal{S}}_{k}\leftarrow\varnothing$
5       while $|{\mathcal{S}}_{k}|<k$ do // Add labels to the prediction set until we hit $k$
@@ 6             $\Delta^{*}\leftarrow-\infty$
                   $\Delta\leftarrow\hat{g}({\mathcal{S}}_{k}\cup\{y\}\,|\,x)-\hat{g}({\mathcal{S}}_{k}\,|\,x)$ // Eval the marginal gain of adding $y$ to ${\mathcal{S}}_{k}$
@@ 9                         $\Delta^{*}\leftarrow\Delta$, $y^{*}\leftarrow y$
            ${\mathcal{S}}_{k}\leftarrow{\mathcal{S}}_{k}\cup\{y^{*}\}$ // Add label offering the largest marginal gain
@@
                   ${\mathcal{S}}\leftarrow{\mathcal{S}}_{k}$ // Update ${\mathcal{S}}$ if ${\mathcal{S}}_{k}$ achieves higher objective value
13
14
15
16return ${\mathcal{S}}$
@@ 5 Experiments with Synthetic Data
Table 1: Empirical average test accuracy achieved by four different (simulated) human experts, each with a different noise value $\gamma$, on their own (None) and using prediction sets constructed with conformal prediction (Naive, Aps, Raps and Saps) and with the greedy algorithm (Greedy) on four synthetic classification tasks. In each classification task, the classifier $f$ used by conformal prediction and the greedy algorithm achieves a different average accuracy $P(Y^{\prime}=Y)$. The number of labels is $L=10$, the size of the calibration set is $m=1000$, and we do not include brute force search because it achieves the same performance as the greedy algorithm. Each cell shows the average and standard deviation over $10$ runs. We denote the best results for each classification task in bold.
@@
Raps $0.341\,\scriptstyle\pm 0.013$ $0.587\,\scriptstyle\pm 0.013$ $0.804\,\scriptstyle\pm 0.014$ $0.941\,\scriptstyle\pm 0.006$
Saps $0.340\,\scriptstyle\pm 0.015$ $0.585\,\scriptstyle\pm 0.012$ $0.804\,\scriptstyle\pm 0.015$ $0.940\,\scriptstyle\pm 0.008$
@@
Raps $0.330\,\scriptstyle\pm 0.012$ $0.566\,\scriptstyle\pm 0.010$ $0.787\,\scriptstyle\pm 0.013$ $0.932\,\scriptstyle\pm 0.008$
Saps $0.329\,\scriptstyle\pm 0.013$ $0.563\,\scriptstyle\pm 0.008$ $0.787\,\scriptstyle\pm 0.014$ $0.932\,\scriptstyle\pm 0.009$
@@
Raps $0.320\,\scriptstyle\pm 0.008$ $0.542\,\scriptstyle\pm 0.012$ $0.760\,\scriptstyle\pm 0.015$ $0.914\,\scriptstyle\pm 0.008$
Saps $0.319\,\scriptstyle\pm 0.009$ $0.534\,\scriptstyle\pm 0.012$ $0.760\,\scriptstyle\pm 0.014$ $0.915\,\scriptstyle\pm 0.009$
@@
Raps $0.316\,\scriptstyle\pm 0.013$ $0.525\,\scriptstyle\pm 0.009$ $0.734\,\scriptstyle\pm 0.015$ $0.896\,\scriptstyle\pm 0.010$
Saps $0.315\,\scriptstyle\pm 0.014$ $0.517\,\scriptstyle\pm 0.011$ $0.734\,\scriptstyle\pm 0.015$ $0.896\,\scriptstyle\pm 0.009$
@@
Experimental setup. We create several synthetic multiclass classification tasks, each with $n=20$ features per sample and varying difficulty. Out of 20 features per sample, only $d=4$ of these features are informative [^11] A feature is informative if its value correlates with the label value. while the rest are drawn at random. Refer to Appendix C for more details about the classification tasks. For each classification task, we generate $19{,}000$ samples, which we split into a training set ($16{,}000$ samples), a calibration set ($1000$ samples), a validation set ($1000$ samples) and a test set ($1000$ samples).
@@
Further, we use the calibration set to calibrate the (softmax) outputs of the logistic regression model $f$ using top-$k$-label calibration [52] with $k=5$. We also use it to estimate the confusion matrices $\bm{C}(\gamma)$ that parameterize the mixture of MNLs used to model (the predictions made by) the simulated human expert, and calculate the quantile $\hat{q}_{\alpha}$ used by conformal prediction. Finally, we use the test set to evaluate the average accuracy achieved by the simulated expert using prediction sets constructed with our greedy algorithm, brute force search and conformal prediction. Here, note that our greedy algorithm and brute force search have access to the true mixtures of MNLs used to model the simulated human expert. In our experiments, we implement conformal prediction using several non-conformity scores:
@@
[…]
where $o(x,y)=|\{y^{\prime}:f_{y^{\prime}}(x)\leq f_{y}(x)\}|$ denotes the ranking of label $y$ according to $f_{y}(x)$ and we decided to omit the randomization for Aps, Raps and Saps as it is only required to achieve exact $1-\alpha$ coverage and it did not have an influence on the empirical average accuracy achieved by the simulated human experts in our experiments. For Raps and Saps, we run the procedure outlined in Appendix E in Angelopoulous et al. [49] to optimize the additional hyperparameters, $k_{regs}$ and $\lambda_{raps}$, for Raps, and $\lambda_{saps}$ for Saps, using the validation set. Further, for each non-conformity score and classification task, we report the results for the $\alpha$ value under which the expert achieves the highest average test accuracy and, to avoid empty sets, we always include the label value with the lowest non-conformity score in the prediction sets. In this context, note that, in practice, one would need to select $\alpha$ using a held-out dataset, however, our evaluation aims to show how our greedy algorithm improves over conformal prediction for any value of $\alpha$. Finally, we repeat each experiment ten times and, each time, we sample different training, calibration, validation and test sets.
@@
Results. We first estimate the average test accuracy achieved by four different (simulated) human experts, each with a different $\gamma$ value, on four classification tasks where the classifier $f$ achieves a different average accuracy $P(Y^{\prime}=Y)$. We report their average test accuracy on their own (None) and when using prediction sets constructed with conformal prediction (Naive, Aps, Raps and Saps), our greedy algorithm (Greedy) and brute force search (Brute Force Search). Table 1 summarizes the results, where we have not included brute force search because it achieves the same performance as our greedy algorithm. The results show that, using the greedy algorithm to construct prediction sets, the experts consistently achieve the highest average accuracy across classification tasks. Moreover, the results also show that, under the prediction sets constructed using the greedy algorithm, the average accuracy achieved by the expert degrades gracefully as $\gamma$ increases whereas, under the prediction sets constructed using conformal prediction, the average accuracy degrades significantly. Refer to Appendix D for additional results for $L\in\{25,50\}$ showing that the relative gain in average accuracy offered by the greedy algorithm increases with the number of labels and noise $\gamma$. Refer to Appendix E for additional results showing that the empirical average coverage achieved by the prediction sets constructed using conformal prediction and our greedy algorithm may be a bad proxy for estimating the average accuracy achieved by human experts using prediction sets.
@@
To better understand why the prediction sets constructed by the greedy algorithm help human experts achieve higher average accuracy than those constructed by conformal prediction, we now look closer into the structure of the prediction sets. Given a ground truth-label $Y=y$, let $\bar{y}=\mathop{\mathrm{argmax}}_{y^{\prime}\neq y}C_{y^{\prime}y}$ be the label that is most frequently mistaken with $y$. Then, we estimate the empirical conditional probability that a prediction set includes $\{y,\bar{y}\}$ given $Y=y$ with the greedy algorithm and conformal prediction. Figure 2 summarizes the results for $\gamma=0.7$ and $P(Y^{\prime}=Y)=0.7$ and $y\in\left\{0,2,6,8\right\}$. Appendix F includes additional results for other configurations. The results show that, with the greedy algorithm, the empirical probability that a prediction set includes $\{y,\bar{y}\}$ given $Y=y$ is much lower (i.e., $2$-$3$x lower) than with conformal prediction despite it creates overall larger prediction sets.
Figure 2: (Left) Confusion matrix $\bm{C}$ for the predictions made by a (simulated) human expert on their own. The label $\bar{y}=\mathop{\mathrm{argmax}}_{y^{\prime}\neq y}C_{y^{\prime}y}$ that is most frequently mistaken with the ground truth-label $y$ is highlighted in red for $y\in\left\{0,2,6,8\right\}$. (Right) Empirical conditional probability that a prediction set includes $\{y,\bar{y}\}$ given $Y=y$ with conformal prediction (Naive, Aps, Raps and Saps) and our greedy algorithm (Greedy). In both panels, $\gamma=0.7$ and $\mathbb{P}(Y^{\prime}=Y)=0.7$.
@@ 6 Experiments with Real Data
In this section, we compare the average accuracy achieved by a simulated human expert using prediction sets constructed with our greedy algorithm, brute force search and conformal prediction on a real multiclass classification task over noisy natural images. The simulated human expert follows the mixture of MNLs introduced in Eq. 4, which is parameterized by the (estimated) confusion matrix of the predictions made by real human experts on their own. [^12] In Appendix G, we evaluate the goodness of fit of the mixture of MNLs to predictions made by real human experts using a support system based on prediction sets [19].
@@
Experimental setup. We experiment with the ImageNet16H dataset [7], which was created using $1{,}200$ natural images from the ImageNet Large Scale Visual Recognition Challenge (ILSRVR) 2012 dataset [55]. More specifically, in the ImageNet16H dataset, each of the above images was used to create four noisy images with different levels of phase noise distortion $\omega\in\{80,95,110,125\}$ and the same ground-truth label $y$ from a label set $\mathcal{Y}$ of size $n=16$. In addition, for each noisy image, the dataset contains (approximately) six label predictions made by human experts on their own. In our experiments, we run and evaluate each method separately by grouping the above noisy images (and expert predictions) according to their level of noise. For each group of images and method, we use the deep neural network classifier VGG-19 [56] after $10$ epochs of fine-tuning as provided by Steyvers et al. [7]. Further, we randomly split the images (and expert predictions) in each group into two disjoint subsets, a calibration set ($800$ images), and a test set ($400$ images). The accuracy of the (pretrained) VGG-19 classifier on the test set is $0.900\pm 0.014$ ($\omega=80$), $0.895\pm 0.009$ ($\omega=95$), $0.857\pm 0.016$ ($\omega=110$) and $0.792\pm 0.016$ ($\omega=125$). We use the calibration set to (i) calibrate the (softmax) outputs of the VGG-19 scores using top-k-label calibration with $k=5$, (ii) estimate the confusion matrix $\bf C$ that parameterizes the mixture of MNLs used to model the simulated human expert, and (iii) calculate the quantile $\hat{q}_{\alpha}$ used by conformal prediction [^13] For Raps and Saps, we further split the calibration set to obtain a (reduced) calibration ($400$ images) and validation ($400$ images) sets to calculate the quantile $\hat{q}_{\alpha}$ and optimize the hyperparameters of Raps and Saps, respectively..We use the test set to evaluate the average accuracy the simulated expert achieves using prediction sets constructed with our greedy algorithm, brute force search and conformal prediction. Here, note that, similarly to in the experiments in synthetic data, our greedy algorithm and brute force search have access to the true mixture of MNLs used to model the simulated human expert. We implement conformal prediction using the same non-conformity scores used in the experiments with synthetic data and, for each non-conformity score and group of images, we report the results for the $\alpha$ value under which the expert achieves the highest average test accuracy. In Appendix H, we report results for all $\alpha$ values. To obtain error bars, we repeat each experiment $10$ times, sampling different calibration and test sets.
@@
Table 2: Average test accuracy achieved by a (simulated) human expert on their own (None), and using the prediction sets constructed with conformal prediction (Naive, Aps, Raps and Saps), and our greedy algorithm (Greedy) on the ImageNet16H dataset. We do not include brute force search because it achieves the same performance as the greedy algorithm. Each cell shows the average and standard deviation over $10$ runs. We denote the best results in bold.
@@
Raps $0.950\,\scriptstyle\pm 0.009$ $0.943\,\scriptstyle\pm 0.010$ $0.914\,\scriptstyle\pm 0.010$ $0.849\,\scriptstyle\pm 0.011$
Saps $0.953\,\scriptstyle\pm 0.010$ $0.942\,\scriptstyle\pm 0.009$ $0.918\,\scriptstyle\pm 0.009$ $0.855\,\scriptstyle\pm 0.007$
@@
None $0.900\,\scriptstyle\pm 0.002$ $0.859\,\scriptstyle\pm 0.003$ $0.771\,\scriptstyle\pm 0.005$ $0.603\,\scriptstyle\pm 0.007$
Results. Table 2 and Figure 3 show the average test accuracy and complementary cumulative distribution (cCDF) of the test per-image accuracy achieved by a simulated human expert using the prediction sets constructed with conformal prediction (Naive, APS, Raps and Saps), our greedy algorithm (Greedy) and brute force search (Brute Force Search) for different values of noise $\omega$. The results show that, similarly as in our experiments with synthetic data, the greedy algorithm achieves the same performance as brute force search. Moreover, they also show that, using greedy algorithm to construct prediction sets, the expert achieves the highest average accuracy in all groups of images except the group with $\omega=80$, where both the greedy algorithm and one of the conformal predictors offer comparable performance. Similarly as in the synthetic experiments, refer to Appendix E for additional results regarding the empirical average coverage achieved by the prediction sets constructed using conformal prediction and our greedy algorithm.
7 Discussion and Limitations
Hardness analysis. In our hardness analysis, our reduction utilizes an instance of our problem in which, for every prediction set, the predictions made by experts follow a mixture of MNLs. As an immediate consequence, this implies that, in general, the problem of finding the prediction set ${\mathcal{S}}$ that maximizes $\hat{g}({\mathcal{S}}\,|\,x)$ is $\mathsf{NP}$-hard to approximate. However, in our experiments, the greedy algorithm is almost always able to find such a set ${\mathcal{S}}$. As a result, we hypothesize that there may be certain conditions on the parameters of the mixture of MNLs under which the problem can be efficiently approximated to a factor less than the size of the label set.
Evaluation. The results of our experiments suggest that the prediction sets constructed using our greedy algorithm may help human experts make more accurate predictions than the prediction sets constructed using conformal prediction. However, one may argue that the difference in performance is partly due to the fact that the non-conformity scores used in conformal prediction do not incorporate information about the distribution of experts’ predictions. Motivated by this observation, it would be important to investigate how to incorporate such information in the definition of non-conformity scores. Moreover, in our experiments, the true distribution of experts’ predictions matches the mixture of MNLs used by the greedy algorithm and brute force search. However, in practice, there may be a mismatch between the true distribution of experts’ predictions and the mixture of MNLs, and this may decrease performance. Finally, in our experiments with real data, the ground truth labels are estimated by aggregating (multiple) predictions by human annotators using majority voting, however, this may introduce additional sources of errors that may influence our results [59].